# Maximum number of points in general position in a random subset of finite $3$-dimensional spaces

### v2

@@ 1 Introduction
For a field $\mathbb{F}$ and a positive integer $d$, let $\mathbb{F}^{d}$ be the $d$-dimensional affine space over $\mathbb{F}$. A point set $S\subseteq\mathbb{F}^{d}$ is in general position if there are no $d+1$ points in $S$ contained in a $(d-1)$-dimensional affine subspace (a $(d-1)$-flat). This notion is closely related to many central problems in Discrete Geometry [8, 9, 11, 5, 16, 3, 23, 18], and we refer the reader to Chapter 8 in [4] for a survey.
In this note, we focus on the case where $\mathbb{F}=\mathbb{F}_{q}$, the finite field with $q$ elements. Let $\alpha(\mathbb{F}_{q}^{d})$ be the maximum possible size of a point set in general position in $\mathbb{F}_{q}^{d}$. Erdős [21] noticed that $\alpha(\mathbb{F}_{q}^{d})\geqslant q$ by the construction $\{(x,x^{2},\ldots,x^{d}):x\in\mathbb{F}_{q}\}$. For the upper bound, it is easy to see that $\alpha(\mathbb{F}_{q}^{d})\leqslant dq$, as $\mathbb{F}_{q}^{d}$ can be partitioned into $q$ disjoint $(d-1)$-flats, and each can contain at most $d$ points. A more careful argument shows that $\alpha(\mathbb{F}_{q}^{d})\leqslant(1+o(1))q$, see for example Lemma 4.1 in [6]. Roche-Newton and Warren [20] started to work on a random version of this problem. Let $\alpha(\mathbb{F}_{q}^{d},p)$ be the maximum possible size of a point set in general position in a $p$-random subset of $\mathbb{F}_{q}^{d}$, where $p$-random means that every point is chosen with probability $p$ independently. They [20] gave essentially tight bounds on $\alpha(\mathbb{F}_{q}^{2},p)$ for $p\in[0,q^{-1}]\cup[q^{-1/3},1]$. Recently, Chen, Liu, Nie, and Zeng [6] determined the order of magnitude of $\alpha(\mathbb{F}_{q}^{2},p)$ up to a polylogarithmic factor for all possible values of $p$ and the order of magnitude of $\alpha(\mathbb{F}_{q}^{d},p)$ when $p$ is in certain ranges.
@@
$\Omega(q^{1/d}/\log q)=\alpha(\mathbb{F}_{q}^{d},p)=O(q^{1/d}\log^{2}q)\quad\textrm{ for }\quad q^{-d+1/d}/\log q\leqslant p<Cq^{-1+1/d}\log^{2}q.$
We will show that 1 follows from the following conjecture about balanced supersaturation. For every set $U\subseteq\mathbb{F}_{q}^{d}$ and every integer $i\in\{1,2,\ldots,d+1\}$, let $\mathcal{G}_{i}(U)$ be the collection of $i$-subsets of $U$ that are not in any $(i-2)$-flat and $\mathcal{S}_{i}(U)\coloneqq\binom{U}{i}\setminus\mathcal{G}_{i}(U)$. For every $r$-uniform hypergraph $\mathcal{H}$ and every set $S\subseteq V(\mathcal{H})$, let $d_{\mathcal{H}}(S)$ be the number of edges in $\mathcal{H}$ containing $S$, and for every integer $i\in\{1,2,\ldots,r\}$, let $\Delta_{i}(\mathcal{H})$ be the maximum value of $d_{\mathcal{H}}(S)$ over all sets $S\subseteq V(\mathcal{H})$ with $|S|=i$.
@@
    $\Delta_{i}(\mathcal{H}_{U})=O(|U|^{d+1-i}/q^{1-(i-1)/d})$ for every integer $i\in\{1,2,\ldots,d\}$,
@@ 2 Balanced supersaturation lemma for $d=3$
In this section, we prove Theorem 4. We remark that the same proof also works for $d=2$, and we will discuss some of the obstacles in higher dimensions in Section 4. We will use the following version of the Chernoff bound, see Corollary A.1.14 in [1].
@@
where $c_{\varepsilon}=\min\{-\ln(e^{\varepsilon}(1+\varepsilon)^{-(1+\varepsilon)},\varepsilon^{2}/2)\}$.
@@
We use the following notation. For every integer $i\in\{1,2,3\}$ and every set $A\in\mathcal{G}_{i}(U)$, let $K_{A}$ be the $(i-1)$-flat containing $A$, $\mathcal{P}_{A}$ be the set of $i$-flats containing $K_{A}$, and
@@
$\mathcal{F}_{A,P}\coloneqq\{F:|F|=4,\,A\subseteq F\subseteq P,\,\textrm{ everysubset of $F$ of size $3$ is in $\mathcal{G}_{3}(U)$}\}$
@@
the pairs $\{a,b\}$ of points in $(P\setminus K)\cap U$ such that $\{z,a,b\}\in\mathcal{S}_{3}(U)$, and $m_{z,K,P}\coloneqq|M_{z,K,P}|$. Let $z_{K,P}$ be the point in $K$ with the largest $m_{z,K,P}$. Note that for $z_{1}\neq z_{2}\in K$, we have $M_{z_{1},K,P}\cap M_{z_{2},K,P}=\varnothing$. Therefore, for every pair $\{x,y\}\subseteq K$ where $z_{K,P}\notin\{x,y\}$, we have
@@
Also, by the assumption that $n_{K}\geqslant 3$, there are
@@
pairs $\{x,y\}\subseteq K\cap U$ such that $z_{K,P}\notin\{x,y\}$, for every $2$-flat $P\supseteq K$. Now, for every $1$-flat $K$ with $3\leqslant n_{K}<n/q^{2/3}$, we have
@@
=(5)${{=}}\Omega\Bigg{(}\sum_{\begin{subarray}{c}P:\textrm{$P$ is a $2$-flat}\\ K\subseteq P\\ n_{P\setminus K}=\Omega(n/q)\end{subarray}}\ \ \sum_{\begin{subarray}{c}\{x,y\}:\{x,y\}\subseteq K\cap U\\ z_{K,P}\notin\{x,y\}\end{subarray}}\binom{n_{P\setminus K}}{2}\cdot\frac{n}{qn_{P\setminus K}}\Bigg{)}$ )
@@
$\mathbb{E}[d_{\mathcal{H}_{A}}(S)]=\mathbb{E}[|\mathcal{H}_{A}|]$ $=\sum_{\begin{subarray}{c}P\in\mathcal{P}_{A}\\ n_{P\setminus K_{A}}=\Omega(n/q)\end{subarray}}\mathbb{E}[|\mathcal{H}_{A,P}|]\leqslant\sum_{\begin{subarray}{c}P\in\mathcal{P}_{A}\\ n_{P\setminus K_{A}}=\Omega(n/q)\end{subarray}}\binom{n_{P\setminus K}}{2}\cdot p_{K,P}$
$=O\Bigg{(}\frac{n}{q}\cdot\sum_{\begin{subarray}{c}P\in\mathcal{P}_{A}\\ n_{P\setminus K_{A}}=\Omega(n/q)\end{subarray}}n_{P\setminus K}\Bigg{)}=O\left(\frac{n^{2}}{q}\right).$
@@
To prove (i), observe that for every $A\in\mathcal{A}$, by Equations 2 and 4, we have
@@
To prove (ii), fix an arbitrary set $S\subseteq U$ with $|S|=i$ ($1\leqslant i\leqslant 3$). First, we consider the case where $i=3$. If $S\in\mathcal{S}_{3}(U)$, then $d_{\mathcal{F}_{A}}(S)>0$ only if $A\subseteq S$; there are at most three such sets $A$. For each of them, we trivially have $d_{\mathcal{F}_{A}}(S)\leqslant n$, implying $\mathbb{E}[d_{\mathcal{H}_{A}}(S)]\leqslant n\cdot p_{A}=O(n/q^{1/3})$ due to our assumption that $n_{K_{A}}\geqslant n/q^{2/3}$ and our definition of $p_{A}$. Hence, we have $\mathbb{E}[d_{\mathcal{H}_{U}}(S)]=O(n/q^{1/3})$. If $S=\{x,y,z\}\in\mathcal{G}_{3}(U)$, then $d_{\mathcal{F}_{A}}(S)>0$ only if $A\subseteq K_{\{x,y\}}$, $K_{\{x,z\}}$ or $K_{\{y,z\}}$. For those sets $A\subseteq K_{\{x,y\}}$, $A$ can be $\{x,y\}$, $\{x,w\}$, or $\{y,w\}$, where $w\in K^{\circ}_{\{x,y\}}$. For all the sets $A$ of the form $\{x,w\}$ or $\{y,w\}$, their contribution to $\mathbb{E}[d_{\mathcal{H}_{U}}(S)]$ is at most
@@
The remaining part of the proof is to prove the actual existence of $\mathcal{H}_{U}\subseteq\mathcal{S}_{4}(U)$ satisfying both (i) and (ii) in all cases. In the discussion above, we proved that in each case, there is a probabilistic construction for $\mathcal{H}_{U}$ such that $\mathbb{E}[|\mathcal{H}_{U}|]=\Omega(n^{4}/q)$ and $\mathbb{E}[d_{\mathcal{H}_{U}}(S)]=O(n^{4-|S|}/q^{1-(|S|-1)/3})$ for every $S\subseteq U$. Then, a simple application of the Chernoff bound guarantees the actual existence of such $\mathcal{H}_{U}$. Indeed, in both Cases 1 and 2, we have $\mathcal{H}_{U}=\cup_{A\in\mathcal{A}}\mathcal{H}_{A}$. Note that for every $A\in\mathcal{A}$ and $S\subseteq U$, we have that $|\mathcal{H}_{A}|$ and $d_{\mathcal{H}_{A}}(S)$ are the sum of mutually independent indicator random variables. This is easy to see in Case 1 and Case 2.2, as $\mathcal{H}_{A}$ is a subset of $\mathcal{F}_{A}$ where each element is kept with a certain probability independently of each other. For Case 2.1, it suffices to notice that $\mathcal{F}_{A,P_{1}}\cap\mathcal{F}_{A,P_{2}}$ for $P_{1}\neq P_{2}\in\mathcal{P}_{A}$ and then recall our definition of $\mathcal{H}_{A}$. Therefore, $\sum_{A\in\mathcal{A}}|\mathcal{H}_{A}|$ and $\sum_{A\in\mathcal{A}}d_{\mathcal{H}_{A}}(S)$ are also sum of mutually independent indicator random variables. We also have
@@ 3 Proof of Claim 3
The proof method used in this section is by now a standard application of the Hypergraph Container Method [2, 22]; we will describe the details to show that the numbers indeed work. We use the following version of the Hypergraph Container Lemma, see Corollary 2.8 in [12] for its proof.
@@
Then, there exists a collection $\mathcal{C}$ of at most $v(\mathcal{G})^{sxv(\mathcal{G})}$ sets of size at most $\left(1-\delta\lambda^{-1}\right)v(\mathcal{G})$ such that for every independent set $I$ in $\mathcal{G}$, there exists some $R\in\mathcal{C}$ with $I\subseteq R$, where $\delta=2^{-s(s+1)}$.
@@
We first prove that for every integer $j\geqslant 0$, there is a collection $\mathcal{C}_{j}$ of at most $(q^{d})^{2(d+1)jq^{1/d}}$ sets of size at most $\max(Tq,(1-\Omega(1))^{j}q^{d})$ such that every set $I\subseteq\mathbb{F}_{q}^{d}$ in general position is in some set from $\mathcal{C}_{j}$. Note that the base case $j=0$ is trivial by taking $\mathcal{C}_{0}=\{\mathbb{F}_{q}^{d}\}$. Given $\mathcal{C}_{j}$, we define $\mathcal{C}_{j+1}$ as follows. For every $U\in\mathcal{C}_{j}$ with $|U|>Tq$, let $\mathcal{H}_{U}$ be the collection given in 2. Note that if $I$ is a subset of $U$ in general position, then it is an independent set in $\mathcal{H}_{U}$. We apply Lemma 8 by letting $\mathcal{G}=\mathcal{H}_{U}$, $\lambda=O(1)$, and $x$ be a real number in $[q^{1/d}/|U|,2q^{1/d}/|U|]$ such that $xv(\mathcal{G})$ and $v(\mathcal{G})/\lambda$ are integers, and obtain a collection $\mathcal{C}_{U}$. Let
@@
$|\mathcal{C}_{j+1}|\leqslant|\mathcal{C}_{j}|\cdot|U|^{(d+1)\cdot\frac{2q^{1/d}}{|U|}\cdot|U|}\leqslant(q^{d})^{2(d+1)jq^{1/d}}\cdot(q^{d})^{2(d+1)\cdot q^{1/d}}=(q^{d})^{2(d+1)(j+1)q^{1/d}}.$
@@
By the assumption that $p=\omega(q^{-d})$ and Lemma 5, we have with high probability that there are at least $pq^{d}/2$ points in a $p$-random subset of $\mathbb{F}_{q}^{d}$. Note that $|\mathcal{S}_{d+1}(\mathbb{F}_{q}^{d})|\leqslant q^{d^{2}+d-1}$, since the number of $(d-1)$-flats in $\mathbb{F}_{q}^{d}$ is $q^{d}$ and every $(d-1)$-flat contains at most $(q^{d-1})^{d+1}=q^{d^{2}-1}$ subsets of size $d+1$. Since $p=o(q^{-d+1/d})$, we have $p^{d+1}q^{d^{2}+d-1}=o(pq^{d})$, so by Markov’s Inequality, we have with high probability that there are at most $pq^{d}/4$ sets in $\mathcal{S}_{d+1}(\mathbb{F}_{q}^{d})$ remaining in the $p$-random subset of $\mathbb{F}_{q}^{d}$, so

### v3

@@ 1 Introduction
For a field $\mathbb{F}$ and a positive integer $d$, let $\mathbb{F}^{d}$ be the $d$-dimensional affine space over $\mathbb{F}$. A point set $S\subseteq\mathbb{F}^{d}$ is in general position if there are no $d+1$ points in $S$ contained in a (d1)(d-1)-dimensional affine subspace (a (d1)(d-1)-flat). This notion is closely related to many central problems in Discrete Geometry [8, 9, 11, 5, 16, 3, 23, 18], and we refer the reader to Chapter 8 in [4] for a survey.
In this note, we focus on the case where $\mathbb{F}=\mathbb{F}_{q}$, the finite field with $q$ elements. Let $\alpha(\mathbb{F}_{q}^{d})$ be the maximum possible size of a point set in general position in $\mathbb{F}_{q}^{d}$. Erdős noticed that $\alpha(\mathbb{F}_{q}^{d})\geqslant q$ by the construction $\{(x,x^{2},\ldots,x^{d}):x\in\mathbb{F}_{q}\}$, see Appendix in [21]. For the upper bound, it is easy to see that $\alpha(\mathbb{F}_{q}^{d})\leqslant dq$, as $\mathbb{F}_{q}^{d}$ can be partitioned into $q$ disjoint (d1)(d-1)-flats, and each can contain at most $d$ points. A more careful argument shows that $\alpha(\mathbb{F}_{q}^{d})\leqslant(1+o(1))q$, see for example Lemma 4.1 in [6]. Roche-Newton and Warren [20] started to work on a random version of this problem. Let $\alpha(\mathbb{F}_{q}^{d},p)$ be the maximum possible size of a point set in general position in a $p$-random subset of $\mathbb{F}_{q}^{d}$, where $p$-random means that every point is chosen with probability $p$ independently. They [20] gave essentially tight bounds on $\alpha(\mathbb{F}_{q}^{2},p)$ for $p\in[0,q^{-1}]\cup[q^{-1/3},1]$. Recently, Chen, Liu, Nie, and Zeng [6] determined the order of magnitude of $\alpha(\mathbb{F}_{q}^{2},p)$ up to a polylogarithmic factor for all possible values of $p$ and the order of magnitude of $\alpha(\mathbb{F}_{q}^{d},p)$ when $p$ is in certain ranges.
@@
$\Omega(q^{1/d}/\log q)=\alpha(\mathbb{F}_{q}^{d},p)=O(q^{1/d}\log^{2}q)\quad\textrm{ for }\quad q^{-d+1/d}/\log q\leqslant p<Cq^{-1+1/d}\log^{2}q,$
where all the implicit constants depend only on $d$.
We will show that 1 follows from the following conjecture about balanced supersaturation. For every set $U\subseteq\mathbb{F}_{q}^{d}$ and every integer $i\in\{1,2,\ldots,d+1\}$, let $\mathcal{G}_{i}(U)$ be the collection of $i$-subsets of $U$ that are not in any (i2)(i-2)-flat and $\mathcal{S}_{i}(U)\coloneqq\binom{U}{i}\setminus\mathcal{G}_{i}(U)$. For every $r$-uniform hypergraph $\mathcal{H}$ and every set $S\subseteq V(\mathcal{H})$, let $d_{\mathcal{H}}(S)$ be the number of edges in $\mathcal{H}$ containing $S$, and for every integer $i\in\{1,2,\ldots,r\}$, let $\Delta_{i}(\mathcal{H})$ be the maximum value of $d_{\mathcal{H}}(S)$ over all sets $S\subseteq V(\mathcal{H})$ with $|S|=i$.
@@
    $\Delta_{i}(\mathcal{H}_{U})=O\left(|U|^{d+1-i}/q^{1-\frac{i-1}{d}}\right)$ for every integer $i\in\{1,2,\ldots,d\}$,
@@ 2 Balanced supersaturation lemma for $d=3$
In this section, we prove Theorem 4. We remark that the same proof also works for $d=2$, and we will discuss some of the obstacles in higher dimensions in Section 4. Our main idea for proving Theorem 4 is to analyze the structure of sets based on at which “dimension” the degeneracy mainly occurs, namely, whether there are many triples in a line in $U$ or not, see Claim 7. In each case, we show the existence of the family $\mathcal{H}_{U}$ satisfying the properties in 2 by constructing it randomly.
We will use the following version of the Chernoff bound, see Corollary A.1.14 in [1].
@@
where $c_{\varepsilon}=\min\{-\ln(e^{\varepsilon}(1+\varepsilon)^{-(1+\varepsilon)}),\,\varepsilon^{2}/2\}$.
@@
We use the following notation. For every integer $i\in\{1,2,3\}$ and every set $A\in\mathcal{G}_{i}(U)$, let $K_{A}$ be the (i1)(i-1)-flat containing $A$, $\mathcal{P}_{A}$ be the set of $i$-flats containing $K_{A}$, and
@@
[…]
@@
the pairs $\{a,b\}$ of points in $(P\setminus K)\cap U$ such that $\{z,a,b\}\in\mathcal{S}_{3}(U)$, and $m_{z,K,P}\coloneqq|M_{z,K,P}|$. Let $z_{K,P}$ be the point in $K\cap U$ with the largest $m_{z,K,P}$. Note that for $z_{1}\neq z_{2}\in K$, we have $M_{z_{1},K,P}\cap M_{z_{2},K,P}=\varnothing$. Therefore, for every pair $\{x,y\}\subseteq K$ where $z_{K,P}\notin\{x,y\}$, we have
@@
Also, by the assumption that $n_{K}\geqslant 3$, we have that for every $2$-flat $P\supseteq K$, there are
@@
pairs $\{x,y\}\subseteq K\cap U$ such that $z_{K,P}\notin\{x,y\}$. Now, for every $1$-flat $K$ with $3\leqslant n_{K}<n/q^{2/3}$, we have
@@
[…]
@@
$\mathbb{E}[d_{\mathcal{H}_{A}}(S)]=\mathbb{E}[|\mathcal{H}_{A}|]$ $=\sum_{\begin{subarray}{c}P\in\mathcal{P}_{A}\\ n_{P\setminus K_{A}}=\Omega(n/q)\end{subarray}}\mathbb{E}[|\mathcal{H}_{A,P}|]\leqslant\sum_{\begin{subarray}{c}P\in\mathcal{P}_{A}\\ n_{P\setminus K_{A}}=\Omega(n/q)\end{subarray}}\binom{n_{P\setminus K_{A}}}{2}\cdot p_{A,P}$
$=O\Bigg(\frac{n}{q}\cdot\sum_{\begin{subarray}{c}P\in\mathcal{P}_{A}\\ n_{P\setminus K_{A}}=\Omega(n/q)\end{subarray}}n_{P\setminus K_{A}}\Bigg)=O\left(\frac{n^{2}}{q}\right).$
@@
To prove (i), observe that for every $A\in\mathcal{A}$, by Equation 2, we have
@@
To prove (ii), fix an arbitrary set $S\subseteq U$ with $|S|=i$ ($1\leqslant i\leqslant 3$). First, we consider the case where $i=3$. If $S\in\mathcal{S}_{3}(U)$, then $d_{\mathcal{F}_{A}}(S)>0$ only if $A\subseteq S$; there are at most three such sets $A$. For each of them, we trivially have $d_{\mathcal{F}_{A}}(S)\leqslant n$, implying $\mathbb{E}[d_{\mathcal{H}_{A}}(S)]\leqslant n\cdot p_{A}=O(n/q^{1/3})$ due to our assumption that $n_{K_{A}}\geqslant n/q^{2/3}$ and our definition of $p_{A}$. Hence, we have $\mathbb{E}[d_{\mathcal{H}_{U}}(S)]=O(n/q^{1/3})$. If $S=\{x,y,z\}\in\mathcal{G}_{3}(U)$, then $d_{\mathcal{F}_{A}}(S)>0$ only if $A\subseteq K_{\{x,y\}}$, $K_{\{x,z\}}$ or $K_{\{y,z\}}$. For those sets $\subseteq K_{\{x,y\}}, AA$ can be $\{x,y\}$, $\{x,w\}$, or $\{y,w\}$, where $w\in K^{\circ}_{\{x,y\}}$. For all the sets $A$ of the form $\{x,w\}$ or $\{y,w\}$, their contribution to $\mathbb{E}[d_{\mathcal{H}_{U}}(S)]$ is at most
@@
The remaining part of the proof is to prove the actual existence of $\mathcal{H}_{U}\subseteq\mathcal{S}_{4}(U)$ satisfying both (i) and (ii) in all cases. In the discussion above, we proved that in each case, there is a probabilistic construction for $\mathcal{H}_{U}$ such that $\mathbb{E}[|\mathcal{H}_{U}|]=\Omega(n^{4}/q)$ and $\mathbb{E}[d_{\mathcal{H}_{U}}(S)]=O(n^{4-|S|}/q^{1-(|S|-1)/3})$ for every $S\subseteq U$. Then, a simple application of the Chernoff bound guarantees the actual existence of such $\mathcal{H}_{U}$. Indeed, in both Cases 1 and 2, we have $\mathcal{H}_{U}=\cup_{A\in\mathcal{A}}\mathcal{H}_{A}$. Note that for every $A\in\mathcal{A}$ and $S\subseteq U$, we have that $|\mathcal{H}_{A}|$ and $d_{\mathcal{H}_{A}}(S)$ are the sum of mutually independent indicator random variables. This is easy to see in Case 1 and Case 2.2, as $\mathcal{H}_{A}$ is a subset of $\mathcal{F}_{A}$ where each element is kept with a certain probability independently of each other. For Case 2.1, it suffices to notice that $\mathcal{F}_{A,P_{1}}\cap\mathcal{F}_{A,P_{2}}=\varnothing$ for $P_{1}\neq P_{2}\in\mathcal{P}_{A}$ and then recall our definition of $\mathcal{H}_{A}$. Therefore, $\sum_{A\in\mathcal{A}}|\mathcal{H}_{A}|$ and $\sum_{A\in\mathcal{A}}d_{\mathcal{H}_{A}}(S)$ are also sum of mutually independent indicator random variables. We also have
@@ 3 Proof of Claim 3
The proof method used in this section is by now a standard application of the Hypergraph Container Method [2, 22]. We use the following version of the Hypergraph Container Lemma, see Corollary 2.8 in [12] for its proof.
@@
Then, there exists a collection $\mathcal{C}$ of at most $v(\mathcal{G})^{s\cdot x\cdot v(\mathcal{G})}$ sets of size at most $\left(1-\delta\lambda^{-1}\right)v(\mathcal{G})$ such that for every independent set $I$ in $\mathcal{G}$, there exists some $R\in\mathcal{C}$ with $I\subseteq R$, where $\delta=2^{-s(s+1)}$.
The main idea for 1 is to apply Lemma 8 to the collections $\mathcal{H}_{U}$ in 2. The property (ii) in 2 guarantees that Equation 12 is always satisfied. We will apply Lemma 8 for a finite number of times and finally show that there is a small collection $\mathcal{C}$ of sets with size at most $T(d)\cdot d$ such that every set in general position in $\mathbb{F}_{q}^{d}$ is contained in some set in $\mathcal{C}$. This enables us to give an upper bound on the number of sets in general position in $\mathbb{F}_{q}^{d}$ and then control the maximum size of it in the random subspace.
@@
We first prove that for every integer $j\geqslant 0$, there is a collection $\mathcal{C}_{j}$ of at most $(q^{d})^{2(d+1)jq^{1/d}}$ sets of size at most $\max(Tq,(1-\Omega(1))^{j}q^{d})$ such that every set $I\subseteq\mathbb{F}_{q}^{d}$ in general position is in some set from $\mathcal{C}_{j}$. Note that the base case $j=0$ is trivial by taking $\mathcal{C}_{0}=\{\mathbb{F}_{q}^{d}\}$. Given $\mathcal{C}_{j}$, we define $\mathcal{C}_{j+1}$ as follows. For every $U\in\mathcal{C}_{j}$ with $|U|>Tq$, let $\mathcal{H}_{U}$ be the collection given in 2. Note that if $I$ is a subset of $U$ in general position, then it is an independent set in $\mathcal{H}_{U}$. We apply Lemma 8 by letting $\mathcal{G}=\mathcal{H}_{U}$, $\lambda=O(1)$, and $x$ be a real number in $q^{1/d}/|U|,2q^{1/d}/|U|]$ such that $xv(\mathcal{G})$ and $v(\mathcal{G})/\lambda$ are integers, and obtain a collection $\mathcal{C}_{U}$. Let
@@
$|\mathcal{C}_{j+1}|\leqslant|\mathcal{C}_{j}|\cdot|U|^{(d+1)\cdot\frac{2q^{1/d}}{|U|}\cdot|U|}\leqslant(q^{d})^{2(d+1)j\cdot q^{1/d}}\cdot(q^{d})^{2(d+1)\cdot q^{1/d}}=(q^{d})^{2(d+1)(j+1)\cdot q^{1/d}}.$
@@
By the assumption that $p=\omega(q^{-d})$ and Lemma 5, we have with high probability that there are at least $pq^{d}/2$ points in a $p$-random subset of $\mathbb{F}_{q}^{d}$. Note that $|\mathcal{S}_{d+1}(\mathbb{F}_{q}^{d})|\leqslant q^{d^{2}+d-1}$, since the number of (d1)(d-1)-flats in $\mathbb{F}_{q}^{d}$ is $q^{d}$ and every (d1)(d-1)-flat contains at most $(q^{d-1})^{d+1}=q^{d^{2}-1}$ subsets of size $d+1$. Since $p=o(q^{-d+1/d})$, we have $p^{d+1}q^{d^{2}+d-1}=o(pq^{d})$, so by Markov’s Inequality, we have with high probability that there are at most $pq^{d}/4$ sets in $\mathcal{S}_{d+1}(\mathbb{F}_{q}^{d})$ remaining in the $p$-random subset of $\mathbb{F}_{q}^{d}$, so